# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ 2.2 The MMT Model
Table 2: Coefficients for the MMT model; $\boldsymbol{\mathrm{A_{2}}}$ is the dissociating molecule and $\boldsymbol{\mathrm{M}}$ is the collision partner
$\mathrm{A_{2}}$ $\mathrm{M}$ $C\,\mathrm{[cm^{3}\,s^{-1}\,mol^{-1}\,K^{-n}]}$ $n$ $T_{D,s}\,\mathrm{[K]}$ $\theta_{v,s}\,\mathrm{[K]}$ $a_{U}$ $U^{*}\,\mathrm{[K]}$
$\mathrm{N_{2}}$ $\mathrm{N_{2}}$ $3.5967\times 10^{18}$ $-0.7017$ 117 529117529117\,529117 529 3 41134113\,4113 411 $0.3868$ 254 556254556254\,556254 556
$\mathrm{N_{2}}$ $\mathrm{N}$ $7.9920\times 10^{17}$ $-0.5625$ 113 957113957113\,957113 957 3 41134113\,4113 411 $0.3668$ 478 708478708478\,708478 708
$\mathrm{N_{2}}$ $\mathrm{O_{2}}$ $5.0420\times 10^{19}$ $-0.9991$ 116 892116892116\,892116 892 3 41534153\,4153 415 $0.3001$ 210 253210253210\,253210 253
$\mathrm{O_{2}}$ $\mathrm{O_{2}}$ $3.6932\times 10^{18}$ $-0.7695$ 60 5406054060\,54060 540 2 28022802\,2802 280 $0.3965$ 57 3435734357\,34357 343
$\mathrm{O_{2}}$ $\mathrm{O}$ $9.2109\times 10^{17}$ $-0.6541$ 60 5526055260\,55260 552 2 28022802\,2802 280 $0.3537$ 237 290237290237\,290237 290
$\mathrm{O_{2}}$ $\mathrm{N_{2}}$ $1.8313\times 10^{15}$ $-0.0223$ 59 3805938059\,38059 380 2 26322632\,2632 263 $0.3620$ 385 466385466385\,466385 466
@@ 2.3 Vibrational relaxation times
In two-temperature nonequilibrium models the relaxation of translational and vibrational modes toward thermal equilibrium is typically described by a source term, as expressed by Eq. (6), in the vibrational energy equation, i.e. Eq. (5). Each species’ contribution is modeled as a Landau-Teller relaxation term, where the numerator represents departure of that species’ vibrational energy $e_{\mathrm{v},s}(T_{\mathrm{v}})$ at local vibrational temperature $T_{\mathrm{v}}$ from the corresponding equilibrium value at the local gas trans-rotational temperature $e_{\mathrm{v},s}(T)$. The denominators in Eq. (6) contain each relaxing diatomic species’ average vibrational relaxation time:
$\langle\tau_{s}^{\mathrm{v}}\rangle=\left(\sum_{q\in S}\frac{x_{q}}{\tau_{s,q}^{\mathrm{v}}}\right)^{-1},\qquad s\in D,$ (18)
@@
itself defined as a weighted average over the inverse of pair-wise relaxation times $\tau_{s,q}^{\mathrm{v}}$, with $x_{q}$ as the collision partner’s mole fraction and $S=\{\mathrm{N_{2}},\mathrm{O_{2}},\mathrm{NO},\mathrm{N},\mathrm{O}\}$ the set of collision partners, depending on the mixture being studied. For air species one would traditionally rely on the Millikan & White correlation [11] to determine these pair-wise relaxation times. However, experimental evidence [32, 33, 34, 35, 36] as well as recent first-principles calculations [25, 26, 30] have revealed significant discrepancies for some species pairings, especially in the high-temperature limit. In this paper we present an alternative expression together with the necessary fit parameters to calculate characteristic vibrational relaxation times for $\mathrm{N_{2}}$, and $\mathrm{O_{2}}$ with their atomic counterparts, which are derived from recent DMS calculations [30] using ab initio PESs. In a subsequent article, additional parameters for a full 5-species air mixture will be presented.
The analytical expression chosen to curve-fit the DMS-derived pair-wise vibrational relaxation times has the form:
$\tau_{s,q}^{\mathrm{v}}=\frac{1}{p}\left(\exp\left(m_{s,q}^{\mathrm{low}}\,T^{-1/3}+n_{s,q}^{\mathrm{low}}\right)+\exp\left(m_{s,q}^{\mathrm{high}}\,T^{-1/3}+n_{s,q}^{\mathrm{high}}\right)\right),\qquad s\in D,q\in S,$ (19)
@@
where $s$ and $q$ are species indices, $p$ is the pressure and $T$ is the trans-rotational temperature. This expression assumes that there are two distinct “low-temperature” and “high-temperature” limiting behaviors for each $\tau_{s,q}^{\mathrm{v}}$, which can be parameterized by separate straight-line segments on a $\log(\tau_{s,q}^{\mathrm{v}}\cdot p)$ vs. $T^{-1/3}$ plot.
Numerical parameters required to evaluate Eq. (19) for individual relaxation pairs are listed in Table 3. For the diatomic species pairs (i.e. $s-q=\mathrm{N_{2}-N_{2}}$ and $\mathrm{O_{2}-O_{2}}$), the DMS predictions tend to agree well with the M&W correlations at lower temperatures, whereas for the diatom-atom pairs the DMS behavior and experiments differ significantly from Millikan & White (for $\mathrm{O_{2}-O}$ in particular see Fig. 7 in Ref. [26]). For all species pairs the high-temperature relaxation behavior predicted by the M&W correlation deviates significantly from that observed in our DMS calculations.
@@
Species $m_{l}$ $n_{l}$ $m_{h}$ $n_{h}$
[…]
@@ 3 Quasi-Classical Trajectory Analysis
This section summarizes how QCT data were used to analyze air dissociation dynamics and correctly scale the reaction rate coefficients as a function of the translational-rotational and vibrational temperatures. Because the trajectory data are constructed with the assumption of a Boltzmann vibrational distribution, non-Boltzmann effects are not considered until later, in Sec. 5. The methods and procedures for generating the QCT data are fully described in Refs. [13, 37].
@@ 3.1 Vibrational Energy Change per Dissociation
[…]
These results indicate the character of molecules that dominate the dissociation process. As described by Bender [13] for the case of $\mathrm{N_{2}+N_{2}}$-dissociation, when conditions are sampled from thermal equilibrium, molecules that dissociate are primarily vibrationally excited. This specificity weakens as temperature increases, because other modes of energy are more able to make up for a lack of vibrational energy. When conditions are sampled with $T_{v}<T$, rotation compensates for vibration and the molecules which tend to dissociate more easily are rotationally excited instead. The present results demonstrate that those findings apply in general for both molecular nitrogen and oxygen dissociation with all the collision partners studied.
@@
[…]
Figure 1: Average change in vibrational energy per dissociation for all 6 dissociation reactions, for their corresponding equilibrium and nonequilibrium test sets. Temperature is normalized by dissociation temperature $T_{D,s}$ and vibrational energy change per dissociation by dissociation energy $D_{0,s}$.
@@ 3.2 Effect of Reactant States
We next consider the effect that each reactant state has on whether dissociation will occur. As described in Ref. [15], the range in probability of dissociation as a function of a specific state quantifies its effect. For example, a low-$v$ molecule is much less likely to dissociate than a high-$v$ molecule, and so the vibrational quantum number has a large effect on dissociation. However, the probability of dissociation varies substantially for different reactions and conditions, which is problematic when aggregating data in the QCT database. Therefore, we divide the probability of dissociation from a specific state by the average probability of dissociation to yield the support factor:
@@
If a reactant state exhibits substantial variation in $\mathcal{S}_{d}$, then that state has a large effect on whether dissociation will occur.
Figure 2 compares the effect that the vibrational and rotational components of the dissociating molecule’s internal energy have on $\mathcal{S}_{d}$. These plots show that the rotational and vibrational energy both affect dissociation, and that the effect of vibrational energy is larger than that of rotational energy. Importantly, these normalized results also collapse, simplifying the development of a model to capture vibrational favoring of dissociation.
(a) Dissociating molecule $\varepsilon_{v}$
(b) Dissociating molecule $\varepsilon_{r}$
[…]

### v2

@@ 2.2 The MMT model formulation
Table 2: Coefficients for the MMT model; $\boldsymbol{\mathrm{A_{2}}}$ is the dissociating molecular species and $\boldsymbol{\mathrm{M}}$ is the collision partner
(a) MMT parameters with variable $T_{\mathrm{D},\mathrm{A_{2}}}$
$\mathrm{A_{2}}$ $\mathrm{M}$ $C\,\mathrm{[cm^{3}\,s^{-1}\,mol^{-1}\,K^{-n}]}$ $n$ $T_{\mathrm{D},\mathrm{A_{2}}}\,\mathrm{[K]}$ $\Theta_{\mathrm{v},\mathrm{A_{2}}}\,\mathrm{[K]}$ $a_{U}$ $U^{*}\,\mathrm{[K]}$ $f_{k}^{\mathrm{NB}}$ $f_{\varepsilon}^{\mathrm{NB}}$
$\mathrm{N_{2}}$ $\mathrm{N_{2}}$ $3.5967\times 10^{18}$ 0.7017-0.7017 117 529117\,529 3 4113\,411 $0.3868$ 254 556254\,556 $0.5$ $0.85$
$\mathrm{N_{2}}$ $\mathrm{N}$ $7.9920\times 10^{17}$ 0.5625-0.5625 113 957113\,957 3 4113\,411 $0.3668$ 478 708478\,708 $0.5$ $0.85$
$\mathrm{N_{2}}$ $\mathrm{O_{2}}$ $5.0420\times 10^{19}$ 0.9991-0.9991 116 892116\,892 3 4153\,415 $0.3001$ 210 253210\,253 $0.5$ $0.85$
$\mathrm{O_{2}}$ $\mathrm{O_{2}}$ $3.6932\times 10^{18}$ 0.7695-0.7695 60 54060\,540 2 2802\,280 $0.3965$ 57 34357\,343 $0.5$ $0.85$
$\mathrm{O_{2}}$ $\mathrm{O}$ $9.2109\times 10^{17}$ 0.6541-0.6541 60 55260\,552 2 2802\,280 $0.3537$ 237 290237\,290 $0.5$ $0.85$
$\mathrm{O_{2}}$ $\mathrm{N_{2}}$ $3.1942\times 10^{17}$ 0.5430-0.5430 62 94962\,949 2 2632\,263 $0.3620$ 385 466385\,466 $0.5$ $0.85$
(b) MMT parameters with $T_{\mathrm{D},\mathrm{A_{2}}}$ held fixed
$\mathrm{A_{2}}$ $\mathrm{M}$ $C\,\mathrm{[cm^{3}\,s^{-1}\,mol^{-1}\,K^{-n}]}$ $n$ $T_{\mathrm{D},\mathrm{A_{2}}}\,\mathrm{[K]}$ $\Theta_{\mathrm{v},\mathrm{A_{2}}}\,\mathrm{[K]}$ $a_{U}$ $U^{*}\,\mathrm{[K]}$ $f_{k}^{\mathrm{NB}}$ $f_{\varepsilon}^{\mathrm{NB}}$
$\mathrm{N_{2}}$ $\mathrm{N_{2}}$ $1.5210\times 10^{17}$ 0.4062-0.4062 113 200113\,200 3 4113\,411 $0.5166$ 365 046365\,046 $0.5$ $0.85$
$\mathrm{N_{2}}$ $\mathrm{N}$ $4.5968\times 10^{17}$ 0.5108-0.5108 113 200113\,200 3 4113\,411 $0.3836$ 503 817503\,817 $0.5$ $0.85$
$\mathrm{N_{2}}$ $\mathrm{O_{2}}$ $3.3947\times 10^{18}$ 0.7470-0.7470 113 200113\,200 3 4153\,415 $0.3664$ 192 401192\,401 $0.5$ $0.85$
$\mathrm{O_{2}}$ $\mathrm{O_{2}}$ $6.9032\times 10^{17}$ 0.6003-0.6003 59 38059\,380 2 2802\,280 $0.4670$ 59 81459\,814 $0.5$ $0.85$
$\mathrm{O_{2}}$ $\mathrm{O}$ $1.6913\times 10^{17}$ 0.4831-0.4831 59 38059\,380 2 2802\,280 $0.3761$ 137 160137\,160 $0.5$ $0.85$
$\mathrm{O_{2}}$ $\mathrm{N_{2}}$ $1.8313\times 10^{15}$ 0.0223-0.0223 59 38059\,380 2 2632\,263 $0.5801$ 174 299-174\,299 $0.5$ $0.85$
@@ 2.3 Vibrational relaxation times
In our proposed two-temperature nonequilibrium model the relaxation of the combined trans-rotational and vibrational modes toward thermal equilibrium is described by Eq. (6), the source term $w_{\mathrm{v}}^{\mathrm{relax}}$ in the vibrational energy balance equation. Each diatomic species’ contribution is modeled as a Landau-Teller relaxation term, where the numerator represents departure of its vibrational energy $e_{\mathrm{v},s}(T_{\mathrm{v}})$ at local vibrational temperature $T_{\mathrm{v}}$ from the corresponding equilibrium value $e_{\mathrm{v},s}(T)$ at the local gas trans-rotational temperature $T$. The denominators in Eq. (6) contain every diatomic species’ average vibrational relaxation time, itself defined as a weighted average over the inverse of pair-wise relaxation times $\tau_{s,q}^{\mathrm{v}}$, with $x_{q}$ as the mole fraction of every collision partner species in the mixture:
$\langle\tau_{s}^{\mathrm{v}}\rangle=\left(\sum_{q\in S}\frac{x_{q}}{\tau_{s,q}^{\mathrm{v}}}\right)^{-1},\qquad s\in D.$ (18)
@@
For the species in partially dissociated air one would traditionally rely on the Millikan and White correlation [11] to determine these pair-wise relaxation times. However, experimental evidence [40, 41, 42, 43, 44], as well as recent DMS studies [31, 32, 37] point to significant departures from the M&W predictions for some species pairings, especially in the high-temperature limit. In this section we present an alternative expression together with the necessary parameters to calculate characteristic vibrational relaxation times for $\mathrm{N_{2}-N_{2}}$, $\mathrm{N_{2}-N}$, $\mathrm{O_{2}-O_{2}}$ and $\mathrm{O_{2}-O}$ pairs. In a subsequent article, we will present additional parameters for a 5-species air mixture. We have derived all these parameters from curve fits to the DMS results of Ref. [37] which relied on the most recent ab initio PESs released by the Minnesota chemists. The analytical expression chosen to curve-fit the DMS-derived relaxation times has the form:
$\tau_{s,q}^{\mathrm{v}}=\frac{c_{0}}{p}\left(\exp\left(m_{s,q}^{\mathrm{low}}\,T^{-1/3}+n_{s,q}^{\mathrm{low}}\right)+\exp\left(m_{s,q}^{\mathrm{high}}\,T^{-1/3}+n_{s,q}^{\mathrm{high}}\right)\right),\qquad s\in D,q\in S,$ (19)
@@
where $s$ and $q$ are species indices, $p$ is the static pressure, factor $c_{0}=1\,\mathrm{atm\cdot s}$ and $T$ is the trans-rotational temperature. This expression assumes that there are two distinct “low-temperature” and “high-temperature” limiting behaviors for each $\tau_{s,q}^{\mathrm{v}}$ which can be parameterized by separate straight-line segments on a $\log(\tau_{s,q}^{\mathrm{v}}\cdot p)$ vs. $T^{-1/3}$ plot.
The DMS data taken from Ref. [37] span a wide temperature range, from an extreme upper limit of $T=100\,000\,\mathrm{K}$ down to values between $2\,000-4\,000\,\mathrm{K}$, depending on the species pair. The precise temperature points we took into account for fitting the high- and low-temperature portions of Eq. (19) respectively were chosen by trial and error and vary slightly for each species pair. Our goal was to obtain an overall curve fit for every individual pair that would reproduce DMS predictions at both ends of the temperature scale, while still adequately bridging between the two limiting straight-line behaviors. We list the resulting numerical parameters for the four relaxation pairs in Table 3 and plot the resulting curve fits for the four collision pairs (blue lines) in Figs. 1(a)-(d), together with the original DMS data points (blue symbols) from Ref. [37]. Each sub-figure also shows the Millikan and White correlation (continuous black lines) and for $\mathrm{N_{2}-N_{2}}$ and $\mathrm{O_{2}-O_{2}}$ pairs the added Park high-temperature correction [45] (dashed black lines).
For the $\mathrm{N_{2}-N_{2}}$ and $\mathrm{O_{2}-O_{2}}$ curve fits in Figs. 1(a) and (c) in particular, the low-temperature behavior is identical to the one predicted by the respective Millikan and White correlation, as we deliberately set their low-temperature parameters in Eq. (19) to be equivalent to the M&W ones. We made this choice, because the M&W correlation for these two relaxing pairs is particularly well-supported by experimental data in the low-temperature limit [11]. At the same time, it is precisely at temperatures below 4 0004\,000 and $3\,000\,\mathrm{K}$ respectively, where the DMS-derived $\tau^{\mathrm{v}}$ data for $\mathrm{N_{2}-N_{2}}$ and $\mathrm{O_{2}-O_{2}}$ is missing. At the upper end of the temperature scale, both species’ DMS curve fits gradually depart from the straight-line Millikan and White trends to predict comparatively longer vibrational relaxation times. This behavior is what motivated Park’s high-temperature correction [45] in the first place. In Figs. 1(a) and (c) we observe qualitative agreement between the DMS curve fits and Park’s correction. However, the correction kicks in at slightly higher temperatures and is more aggressive than the DMS predictions in raising the vibrational relaxation times above those of the M&W correlation.
For the $\mathrm{N_{2}-N}$ pair, across most of the temperature range shown in Fig. 1(b), the DMS calculations predict vibrational relaxation times somewhat shorter than those obtained with Millikan and White. However, as already reported in Ref. [32], we observe the widest gap between Millikan and White and DMS for the $\mathrm{O_{2}-O}$-pair. We show its DMS curve fit in Fig. 1(d) and see it predicting a nearly constant value on a $\log p\cdot\tau^{\mathrm{v}}$ vs. $T^{-1/3}$ plot. For these diatom-atom pairs the vibrational relaxation times are particularly difficult to access experimentally, as they only become significant in partially dissociated mixtures. Thus, shock-tube experiments attempting to measure them have to operate at higher post-shock temperatures, precisely where the M&W correlation is unlikely to provide a good fit to the data. Therefore, we relied exclusively on the available DMS data and made no attempt to force any part of either pair’s curve fit to agree with prior M&W correlation predictions.
@@
Species $m^{\mathrm{low}}$ $n^{\mathrm{low}}$ $m^{\mathrm{high}}$ $n^{\mathrm{high}}$
pair sqs-q $\mathrm{[K^{1/3}]}$ $\mathrm{[-]}$ $\mathrm{[K^{1/3}]}$ $\mathrm{[-]}$
@@
Figure 1: Vibrational relaxation times for nitrogen and oxygen. DMS data [37] and curve fits (blue), M&W correlation [11] plus Park’s high-temperature correction [45] (black).
@@ 3 Quasi-classical trajectory analysis
In this section we summarize how we used an extensive QCT database to analyze air dissociation dynamics and correctly scale the reaction rate coefficients as a function of the translational-rotational and vibrational temperatures. This summary constitutes a highly condensed version of the analysis carried out in chapter 3 of Ref. [25]. Because we constructed the trajectory data with the assumption of a Boltzmann vibrational distribution, we do not consider non-Boltzmann effects until later, in Sec. 5. The methods and procedures for actually generating the QCT data are fully described in Refs. [15, 39].
@@ 3.1 Vibrational energy change per dissociation
We first present QCT-derived data for the average vibrational energy removed, or alternatively the negative energy change per dissociation [^6] For all dissociation reactions considered in this work, $\nu_{s,\mathrm{diss}}^{b}-\nu_{s,\mathrm{diss}}^{f}=-1$, which means that vibrational energy change and energy removed are always related as: $\langle\Delta\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}=\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$. As discussed at the end of Sec. 2.1, this variable quantifies the effect that dissociation has on the vibrational energy content of the gas. The raw QCT data for $\mathrm{O_{2}\,by\,M}$ dissociation with collision partners $\mathrm{M}=\{\mathrm{O_{2},O,N_{2}}\}$ and $\mathrm{N_{2}\,by\,M}$ dissociation with partners $\mathrm{M}=\{\mathrm{N_{2},N,O_{2}}\}$ is shown in Fig. 3.22(a) of Ref. [25] for equilibrium conditions ($T\!=\!T_{\mathrm{v}}$). Oxygen dissociation results were gathered within the range $4\,000\,\mathrm{K}\leq T\leq 13\,000\,\mathrm{K}$, whereas nitrogen results are available for the interval $8\,000\,\mathrm{K}\leq T\leq 30\,000\,\mathrm{K}$. For the same collision pairs Fig. 3.22(b) of Ref. [25] presents QCT-derived vibrational energy change per dissociation at nonequilibrium conditions characterized by a range of $(T,T_{\mathrm{v}})$ combinations. For oxygen dissociation the sweep over $T_{\mathrm{v}}$ is shown for constant $T=10\,000\,\mathrm{K}$, whereas the nitrogen dissociation sweep is shown for a trans-rotational temperature of $20\,000\,\mathrm{K}$. We normalize this change in vibrational energy per dissociation by the corresponding dissociation energy $D_{0,s}$, and the temperature by the corresponding dissociation temperature, $T_{\mathrm{D},s}$. We plot this normalized vibrational energy change per collision in Fig. 2, where we can see that the normalized vibrational energy change collapses to within about 10% when normalized by the dissociation energy for both vibrational equilibrium and nonequilibrium conditions.
These results indicate the character of molecules that dominate the dissociation process. As described by Bender [15] for the case of $\mathrm{N_{2}+N_{2}}$ dissociation, when conditions are sampled from thermal equilibrium, molecules that dissociate are primarily vibrationally excited. This specificity weakens as temperature increases, because other modes of energy are more able to make up for a lack of vibrational energy. When conditions are sampled with $T_{\mathrm{v}}\!<\!T$, rotation compensates for vibration and the molecules which tend to dissociate more easily are rotationally excited instead. The present results demonstrate that those findings apply in general for both molecular nitrogen and oxygen dissociation with all the collision partners studied. Refer to Sec. 3.4 of Ref. [25] for further details.
@@
(b) Nonequilibrium test set ($T=20\,000\,\mathrm{K}$ for $\mathrm{N_{2}\,by\,M}$ and $T=10\,000\,\mathrm{K}$ for $\mathrm{O_{2}\,by\,M}$).
Figure 2: Average change in vibrational energy per dissociation normalized by $D_{0,s}$ for all 6 dissociation reactions, temperatures normalized by $T_{\mathrm{D},s}$.
@@ 3.2 Effect of reactant states
We next consider the effect that each reactant state has on whether dissociation will occur. As described in Ref. [22], the range in probability of dissociation as a function of a specific state quantifies its effect. For example, a low-$v$ molecule is much less likely to dissociate than a high-$v$ molecule and thus the vibrational quantum number has a large effect on dissociation. However, the probability of dissociation varies substantially for different reactions and conditions, which is problematic when aggregating data in the QCT database.
In order to facilitate the analysis, for every reactant pair of interest (either $\mathrm{O_{2}}(v)\,\mathrm{by}\,\mathrm{M}$, or $\mathrm{N_{2}}(v)\,\mathrm{by}\,\mathrm{M}$) we may compute the vibration-specific, thermally averaged dissociation probability $P_{d}\,(v)=P_{d}\,(\varepsilon_{s}(v),T)$ as the ratio of all quasiclassical trajectories where the dissociating molecule initially possesses vibrational energy $\varepsilon_{s}(v)$ relative to the total number of trajectories for the reactant pair, regardless of outcome. We further calculate the Boltzmann-weighted, thermally averaged dissociation probability $P_{d}=P_{d}(T,T_{\mathrm{v}})=\sum_{v=0}^{n_{v,s}-1}\{P_{d}\,(v)\exp\left(-\varepsilon_{s}(v)/\mathrm{k_{B}}T_{\mathrm{v}}\right)\}/Q_{\mathrm{v},s}(T_{\mathrm{v}})$. The ratio of both quantities yields the support factor:
@@
If a reactant state exhibits substantial variation in $\mathcal{S}_{d}$, then that state has a large effect on whether dissociation will occur. This is illustrated by Fig. 3.25(a) of Ref. [25] for the special case of $\mathrm{O_{2}}(v)+\mathrm{N_{2}}$ dissociation at $T\!=\!T_{\mathrm{v}}=10\,000\,\mathrm{K}$. At these conditions the support factor is seen to increase over nearly 4 orders of magnitude from the lowest to the highest-lying $v$-states of oxygen. When the support factor is then plotted for a range of temperatures, as in Fig. 3.26 of Ref. [25], it also becomes clear that the selectivity for dissociation on $v$-level is highest at low $T$ and gradually becomes less pronounced at high temperatures.
The concept of the support for dissociation may be generalized. Instead of plotting against discrete $v$-levels, the support factor may alternatively be parameterized in terms of the rotational, vibrational or overall internal energies, or the relative translational energy of the collision pair. This analysis is done in Sec. 3.5.2 of Ref. [25] for all six oxygen and nitrogen dissociation reactions at $T\!=\!T_{\mathrm{v}}=10\,000\,\mathrm{K}$. Although the support is seen to depend on all these quantities to some extent, the strongest and most direct dependence is observed for vibrational and rotational energies of the dissociating diatom. Therefore, in Fig. 3 we compare the effect that the vibrational and rotational components of the dissociating molecule’s internal energy have on $\mathcal{S}_{d}$. These plots show that the rotational and vibrational energy both affect dissociation, and that the effect of vibrational energy is larger than that of rotational energy. Importantly, these normalized results also collapse, simplifying the development of a model to capture vibrational favoring of dissociation.
(a) Dependence on vibrational energy
(b) Dependence on rotational energy
Figure 3: Support for single/simple dissociation for all dissociation reactions, at $T=T_{\mathrm{v}}=10\,000\,\mathrm{K}$.